OpenStudy (anonymous):

Given g(x)= 6x^2/ sqrt(4+3x^2) find and simplify

OpenStudy (anonymous):

\[g(\frac{ 2 }{ \sqrt{3} }\tan \theta)\] where \[\theta \epsilon (\frac{ \pi }{ 2 }, \pi)\]

OpenStudy (tkhunny):

"Find"?

OpenStudy (anonymous):

And simplify

OpenStudy (tkhunny):

So, \(\tan(\pi) = \;??\;and \tan(\pi /2) = \;??\)

OpenStudy (anonymous):

Not sure

OpenStudy (tkhunny):

Then you cannot solve this problem. Be sure and we can continue. \(\tan(\theta) = \dfrac{\sin(\theta)}{\cos(\theta)}\)